# FERGI: Automatic Annotation of User Preferences for Text-to-Image Generation from Spontaneous Facial Expression Reaction

### v2

@@ 6.2.3 Baseline Scores
To be consistent with the evaluation of AUcomb valence score, a leave-one-participant-out (LOPO) procedure was applied in the fitting of the three weights $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$; we used grid search with a granularity of 0.1 to optimize the accuracy of image preference prediction. Since the scales of the scores are different, they were each z-scored using their mean and standard deviation on the training set. We obtained the same results of $w_{\mathrm{IR}}=0.1$, $w_{\mathrm{Pick}}=0.6$, and $w_{\mathrm{HPSv2}}=0.3$ for all LOPO training sets.
6.2.4 AUcomb Valence Score: Independent Preference Prediction for Selected Image Pairs
We first evaluate the AUcomb valence score on its ability to independently predict user preferences between pairs of images generated by the same text prompt. Ground truth labels of preferences between images are derived from the image ranking at the end of each session, which resulted in user preference data on a total of 3534 image pairs from the 30 included participants.
Considering the randomness of AU activation and the imperfection of the AU estimation model, a small difference between two AUcomb valence scores may not be informative for inferring preference between the two associated images. Therefore, we apply a threshold $d$ for making predictions: only when the difference between the AUcomb valence scores of two images is larger than or equal to the threshold $d$, do we predict that the user prefers the image with the higher AUcomb valence score; otherwise, we abstain from making a prediction.
There are 3 parameters to fit for the proposed algorithm: the exponential decay coefficient $k$ (as in Eqs. 3 and 4), the weight for AU4 $w_{\mathrm{AU4}}$ (as in Eq. 5; $w_{\mathrm{AU12}}$ is then automatically set by Eq. 6) and the difference threshold $d$. We fit these three parameters with an LOPO procedure using grid search (from 0.1 to 2.0 with a granularity of 0.1 for $k$, from 0.0 to 1.0 with a granularity of 0.1 for $w_{\mathrm{AU4}}$, and from 0.00 to 1.00 with a granularity of 0.02 for $d$) to optimize
$n_{\mathrm{correct}}-2n_{\mathrm{incorrect}},$ (9)
where $n_{\mathrm{correct}}$ and $n_{\mathrm{incorrect}}$ represents the number of correct predictions and incorrect predictions on image pair preference.
Table 1: Accuracy of image preference prediction on the 19.38% of all data (685 out of 3534 pairs) as selected by the AUcomb valence model. (Image pairs with AUcomb valence score difference below the trained difference threshold are not included.)
Model Acc.
CLIP Score 58.25
Aesthetic Score 51.39
BLIP Score 49.78
ImageReward Score 62.19
PickScore 61.31
HPS V2 Score 62.04
Ensemble Baseline Score 66.57
AUcomb Valence Score 72.70
The results are shown in Tab. 1. Based on the trained difference thresholds, the AUcomb valence model makes predictions on 19.38% of the image preference data (685 out of 3534 pairs). For the selected image pairs, our AUcomb valence score has the best accuracy of 72.70%, outperforming the ensemble of the three pre-trained human preference scoring models by a margin of 6.13%.
Our results have high stability. The parameter values fit in the LOPO procedure were generally consistent. We obtained $k=0.5$, $w_{\mathrm{AU4}}=0.5$, $w_{\mathrm{AU12}}=0.5$, and $d=0.26$ for 20 out of the 30 total training sets.
@@ 6.2.4 AUcomb Valence Score: Independent Preference Prediction for Selected Image Pairs
6.2.5 Integrate AUcomb Valence Score with Pre-trained Scoring Models
AUcomb valence score has a great performance of 72.70% on the subset of the image pairs with distinguishable AUcomb valence scores, but if we force it to make a prediction on every image pair in the dataset, its accuracy falls to 58.23%, still significantly better than chance (two-sided binomial test: $p<5\times 10^{-22}$) but lower than that of the ImageReward score, PickScore, HPS v2 score, and ensemble baseline score (see Tab. 2).
However, besides making independent predictions, the AUcomb valence score can also be integrated with pre-trained scoring models to enhance their consistency with human preferences, which will be applicable to all generated images with valid facial expression reaction data. We integrate the pre-trained scoring models with AUcomb valence score with a simple linear combination:
$s_{m}+a_{m}s_{\mathrm{AUcomb}},$ (10)
@@ 6.2.5 Integrate AUcomb Valence Score with Pre-trained Scoring Models
where $m$ represents the specific pre-trained scoring model, $s_{m}$ the score from model $m$, and $a_{m}$ the weight for integrating $m$ and the AUcomb valence score.
@@
We separately investigated the effect of integrating the AUcomb valence score with each of the three scoring models trained on large human preference datasets (ImageReward, PickScore, and HPS v2), and their ensemble. Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 10.0 with a granularity of 0.1); for the weights of the ensemble baseline score (Eq. 7), exponential decay coefficient for the AU4 and AU12 valence scores (Eqs. 3 and 4), and the weighting between AU4 and AU12, we directly used the values fit previously.
@@
Table 2: Accuracy of image preference prediction on all data. The second column shows the accuracy of the models given in the first column when they make predictions independently while the third column shows the accuracy of the best few when they are integrated with the AUcomb valence score.
@@
CLIP Score 57.50 /
Aesthetic Score 49.92 /
BLIP Score 49.92 /
AUcomb Valence Score 58.23 /
ImageReward Score 60.44 63.67
PickScore 63.36 65.42
HPS V2 Score 61.94 65.53
Ensemble Baseline Score 65.79 67.49
The results are shown in Tab. 2. Integration with AUcomb valence score improves the performance of all of the four baseline models. Specifically, by further integrating all three pre-trained human preference scoring models together with the AUcomb valence score, we achieve the highest accuracy of 67.49%, outperforming the ensemble of the three models by a margin of 1.70%.
@@ 7 Discussion and Conclusion
We propose that annotation of user preferences for text-to-image generation can be automated with analysis of user facial expression reaction to generated images, which can potentially increase the efficiency and scalability of human preference data collection. We present the FERGI dataset, consisting of video recordings of facial expression reaction to text-to-image generation, and show that multiple AUs are correlated with the participant’s evaluation of and emotional reaction to the generated image.
Specifically, we developed an AUcomb valence score to represent the user’s evaluation of the image reflected by AU4 and AU12 activations, which is useful in automatically annotating user preferences between pairs of images generated from the same input text prompt.
Firstly, the AUcomb valence score independently can select a subset of available data with large score difference and achieve a high accuracy in preference annotation of the selected data, substantially outperforming state-of-the-art pre-trained scoring models, including their ensemble. Therefore, it can potentially be used to selectively annotate large-scale user preference data used for fine-tuning scoring models. Note that automatic annotation will enable collection of vastly more human preference data from real users, so even if only a small portion are usable, its scale will still be much larger than manually annotated datasets.
Secondly, the AUcomb valence score can be integrated with pre-trained scoring models to improve their accuracy in image preference prediction, which can be potentially helpful when the scoring models are used to label human preferences for fine-tuning text-to-image generative models. Importantly, previous work has shown that even if the accuracy of the predictions is not very high, it can still be effective for fine-tuning of text-to-image models: Xu et al. [75] used the reward model ImageReward (trained from manually labeled human preferences), which only has an accuracy of 65.14% on their own dataset and 60.44% on our FERGI dataset, to fine-tune SD v1.4 and obtained a new model whose generated images were preferred by humans 58.79% of the time over those generated by the original SD v1.4. (see Tab. 4 and Fig. 6 of [75]).
Thirdly, our AUcomb valence score may be useful for understanding individual user preferences to create personalized image generation models.
In conclusion, we have demonstrated the feasibility of automatically annotating user preferences for image generation from facial expression reaction. The application of this method is not limited to text-to-image generation but may also be potentially applicable to other image generation tasks, such as image-to-image translation [53], image inpainting [21], and super-resolution [70].
@@ 8 Acknowledgments
A Data Collection
This section provides more details of the collection of the FERGI dataset not provided in Sec. 3. Figs. S1, S2, S3, S4, S5, S6 and S7 display screenshots of an example of one session in data collection.
A.1 Participants
39 participants [^1] The user IDs are up to 040 as shown in Tab. S1 because the same participant used both 027 and 028 due to some technical issues. Though, no data was collected under the user ID 027. were recruited from the SONA system of University of California San Diego (UCSD). They underwent the informed consent process and received instruction for the study from the researcher via a videoconference and completed the study asynchronously afterwards on their personal computers. The videos of 3 participants failed to record and displayed only a logo of OBS Studio (the reason has not yet been fully understood by the researchers); 1 participant dropped out of study under the instruction of the researcher because the only available personal device was a tablet, which is not compatible with some procedures of the study; 1 participant did not complete any study sessions asynchronously at all after receiving the instruction in the videoconference. Among the other 34 participants with valid data and videos recorded, only 1 participant chose not to permit video recordings being shared with other researchers for research purposes; therefore, we decided to also exclude the data of this participant in our analysis for enhancing reproducibility of our study within the research community. We ended up with a dataset of 33 participants that will be made available to researchers for research purposes.
A.2 Prompt Creation
@@ A.2.1 Structured Input
After participants finish with the input sections and click “generate prompt”, a prompt will be generated based on all of the inputs with LLM (OpenAI gpt3.5-turbo model). Participants can freely adjust the generated prompt as the final prompt. Screenshots of an example of the inputs and prompt generation for a structured input session are shown in Figs. S1 and S2.
A.2.2 Free-Form Input
@@
A.3 Image Presentation
Screenshots of an example of the pages for image presentation and its baseline are shown in Figs. S4 and S5.
A.3.1 Image Annotation
@@ B AU Model Training
B.1 Datasets
The DISFA dataset [50] contains facial video recordings of 27 participants’ spontaneous facial expression while viewing video clips with approximately $130,000$ frames in total. Each frame is manually annotated by a human expert with intensities of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop) on a scale of 0 to 5222See https://melindaozel.com/facs-cheat-sheet/ for a visual reference guide for FACS.
@@ B.3 Model Training
For model training, we employ the Adam optimizer with an initial learning rate of $10^{-4}$ for parameters of the last layer and an $10^{-5}$ for other parameters, a weight decay of $5\times 10^{-4}$, and a batch size of 64. We train the model on all data from the DISFA and DISFA+ datasets [50, 48] for a total of 3 epochs using a single NVIDIA GeForce GTX 1080Ti 11G GPU.
C Data Filtering

### v3

@@ 6.2.1 Pre-trained Scoring Models
A leave-one-participant-out (LOPO) procedure was applied in the fitting of the three weights $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$; we used grid search with a granularity of 0.1 to optimize the accuracy of image preference prediction. Since the scales of the scores are different, they were each z-scored using their mean and standard deviation on the training set. We obtained the same results of $w_{\mathrm{IR}}=0.1$, $w_{\mathrm{Pick}}=0.6$, and $w_{\mathrm{HPSv2}}=0.3$ for all LOPO training sets.
6.2.2 FAU-Net Valence Score
We aim to use the estimated AU activation values to automatically score user preferences between images generated from the same input text prompt and propose that this scoring more directly reflects the user’s reaction to the generated image and is complementary to the pre-trained scoring models. To achieve this goal, we train an FAU-Net (Facial Action Units Neural Network) to do automatic scoring. In the FAU-Net, the estimated activation values of the 12 investigated AUs are used as the inputs; each activation value is first preprocessed with a linear transformation with trained parameters and sigmoid function ($\sigma(a\cdot\alpha+b)$ with the initializations of $a=1$ and $b=-1$ for encouraging transformations of positive correlation for better interpretability); then the preprocessed AU activation values are used as inputs to a hidden layer with 16 neurons which feeds into a single ouput neuron representing the AU evaluation score of the image; this output unit receives inputs from both the 16 neurons of the hidden layer and the 12 preprocessed AU activation values. The sigmoid function is used as the activation function for all neurons. The network is trained with a RankNet loss function using the reaction clips and ranking order given by the participant for image pairs generated from the same text prompts and an L1 regularization with a coefficient of $10^{-4}$. For model training, we use the Adam optimizer with an initial learning rate of $10^{-3}$ and a single batch for all data to train the neural network for 300 epochs using a single NVIDIA RTX A6000 48GB GDDR6 GPU for about 15 minutes.
Our FAU-Net is trained with an LOPO procedure and outputs an FAU-Net valence score $s_{\mathrm{FAU}}$ reflecting the evaluation of the generated image derived from the facial expression reaction. Note that the AU estimation model is frozen and not trained with the FERGI data.
6.2.3 Integrate FAU-Net Valence Score with Pre-trained Scoring Models
@@
We integrate our FAU-Net valence score and the pre-trained scoring model with a simple linear combination:
$s_{m}+a_{m}s_{\mathrm{FAU}},$ (5)
@@
where $m$ represents the specific pre-trained scoring model, $s_{m}$ the score from model $m$, and $a_{m}$ the weight for integrating $m$ and the FAU-Net valence score.
@@
We separately investigated the effect of integrating the FAU-Net valence score with each of the three scoring models trained on large human preference datasets (ImageReward, PickScore, and HPS v2), and their ensemble. Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 10.0 with a granularity of 0.1); the FAU-Net weights and the weightings $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, $w_{\mathrm{HPSv2}}$ computed in Eq. 3 are not modified at this stage.
@@
Table 2: Accuracy of image preference prediction for a total of 3734 image pairs. The second column shows the accuracy of the models given in the first column when they make predictions independently while the third column shows the accuracy of the best few when they are integrated with the FAU-Net valence score.
@@
CLIP Score 57.74 /
Aesthetic Score 50.32 /
BLIP Score 50.05 /
FAU-Net Valence Score 59.35 /
ImageReward Score 60.47 65.21
PickScore 63.28 66.28
HPS V2 Score 62.00 66.87
Ensemble Baseline Score 65.80 68.64
(a) Selected by
ImageReward score
(b) Selected by
PickScore
(c) Selected by
HPS V2 Score
(d) Selected by
FAU-Net Valence Score
Figure 4: FAU-Net valence score is complementary to the other pre-trained scoring models. Each subfigure shows how annotation accuracy changes if only a subset of the data are selected to annotate by the different scoring models (given in the subcaption label), with the x-axis representing the proportion of data selected (image pairs are selected based on the largest absolute difference of scores for the two images) and the y-axis representing the annotation accuracy within the selected subset. Each model is good at estimating the image pairs that it will perform well on as evidenced by increased accuracy for a smaller selected proportion. Note the blue, green, and orange curves are almost on top of each other.
(a) ${\color[rgb]{1,0,0}w_{h(1)-o}\approx-0.58}$
(b) ${\color[rgb]{1,0,0}w_{h(2)-o}\approx-0.53}$
(c) ${\color[rgb]{1,0,0}w_{h(3)-o}\approx-0.50}$
(d) ${\color[rgb]{1,0,0}w_{h(4)-o}\approx-0.49}$
(e) ${\color[rgb]{1,0,0}w_{h(5)-o}\approx-0.46}$
(f) ${\color[rgb]{1,0,0}w_{h(6)-o}\approx-0.45}$
(g) ${\color[rgb]{1,0,0}w_{h(7)-o}\approx-0.45}$
(h) ${\color[rgb]{1,0,0}w_{h(8)-o}\approx-0.42}$
(i) ${\color[rgb]{1,0,0}w_{h(9)-o}\approx-0.37}$
(j) ${\color[rgb]{1,0,0}w_{h(10)-o}\approx-0.22}$
(k) ${\color[rgb]{1,0,0}w_{h(11)-o}\approx-0.13}$
(l) ${\color[rgb]{1,0,0}w_{h(12)-o}\approx-0.12}$
(m) ${\color[rgb]{0.0,0.7,0.0}w_{h(13)-o}\approx 0.30}$
(n) ${\color[rgb]{0.0,0.7,0.0}w_{h(14)-o}\approx 0.31}$
(o) ${\color[rgb]{0.0,0.7,0.0}w_{h(15)-o}\approx 0.42}$
(p) ${\color[rgb]{0.0,0.7,0.0}w_{h(16)-o}\approx 0.56}$
Figure 5: Weights of hidden nodes. Each subfigure shows the weights from the input preprocessed activation values of the 12 AUs to a hidden node in the FAU-Net, with each subcaption illustrating the weight from the hidden node to the output node (red/green for negative/positive weights). The hidden nodes are ordered by the weights from them to the output node in this figure.
Figure 6: Weights of the output node. The weights from the input preprocessed activation values of the 12 AUs to the output node in the FAU-Net.
The results are shown in Tab. 2. Integration with FAU-Net valence score improves the performance of all four baseline models. Specifically, by further integrating all three pre-trained human preference scoring models together with the FAU-Net valence score, we achieve the highest accuracy of 68.64%, outperforming the ensemble of the three models by a margin of 2.84% ($p<0.01$ with a two-proportion z-test).
6.2.4 Complementarity between FAU-Net and Pre-trained Scoring Models
The complementarity between FAU-Net and pre-trained scoring models is demonstrated in Fig. 4. The figure shows how annotation accuracy changes if only a subset of the data are selected to annotate by the different scoring models (image pairs are selected based on the largest absolute difference scores for the two images given by the associated scoring model). We can clearly see that the three pre-trained scoring models are highly correlated showing similar performance curves (almost overlapping). In contrast, our FAU-Net valence score selects different image pairs, explaining why the FAU-Net valence scores combine synergistically with the other scores.
6.2.5 FAU-Net Analysis
We analyze the weights of the trained FAU-Net to better understand the model. Figure 5 shows the weights from the input preprocessed AU activation values to the nodes in the hidden layer (given that the preprocessings of activation values are still all increasing functions after model training) with the subcaptions illustrating the weight from the corresponding hidden node to the output node. Generally, for hidden nodes positively affecting the output score, they receive positive weights from the activations of AU2, AU6, AU12, AU17, AU20, and AU26 and negative weights from the activations of AU4, AU5, AU9, AU17, and AU25, and for hidden nodes negatively affecting the output score, they receive weights of the opposite signs from the AUs. Similar trends are observed for the weights from the input preprocessed AU activation values and the output node as shown in Fig. 6.
@@ 7 Discussion and Conclusion
We propose that user facial expression reactions to text-to-image generation can be used to score user preferences for the generated images, which is complementary to existing scoring models which estimate human preferences based only on the input text prompts and generated images. We present the FERGI dataset, consisting of video recordings of facial expression reaction to text-to-image generation, and show that multiple AUs are correlated with the participant’s evaluation of and emotional reaction to the generated image.
Specifically, we developed an FAU-Net that receives the detected AU activations in the facial expression reaction as inputs and outputs an FAU-Net valence score as the estimation of the user preference. Integrating the FAU-Net valence score with pre-trained scoring models improves their accuracy in image preference prediction, which can be potentially helpful when the scoring models are used to label human preferences for fine-tuning text-to-image generative models.
Importantly, previous work has shown that even if the accuracy of the predictions is not very high, it can still be effective for fine-tuning of text-to-image models: Xu et al. [79] used the reward model ImageReward (trained from manually labeled human preferences), which only has an accuracy of 65.14% on their own dataset and 60.47% on our FERGI dataset, to fine-tune SD v1.4 and obtained a new model whose generated images were preferred by humans 58.79% of the time over those generated by the original SD v1.4. (see Tab. 4 and Fig. 6 of [79]).
We also demonstrated the complementarity between our FAU-Net and the pre-trained scoring models, explaining why integrating them effectively improves the annotation accuracy. As they are respectively better at annotating different sets of image pairs, it is also potentially feasible to use them to fine-tune each other on data where one model gives larger absolute difference of scores for the two images.
In conclusion, we have demonstrated the feasibility of automatically scoring user preferences for image generation from facial expression reaction. The application of this method is not limited to text-to-image generation but may also be potentially applicable to other image generation tasks, such as image-to-image translation [57], image inpainting [22], and super-resolution [74].
@@ 8 Acknowledgments
A Supplementary Visuals
To better illustrate the idea of our paper, we present Fig. S1 showing an example of a strong AU4 activation in response to a low-quality image generation and Fig. S2 an example of a strong AU12 activation in response to a high-quality image generation.
Figure S1: An example of AU4 activation in response to a low-quality image generation. (a) The input text prompt. (b) The facial expression of the participant before seeing the generated image. (c) The generated image. (d) The facial expression of the participant after seeing the generated image (the frame with the highest estimated AU4 intensity). In the annotation survey, the participant gave an overall rating of 2, an image-text alignment rating of 2, and a fidelity rating of 1; she identified the image with the issues “output contains unwanted content that was not mentioned in the text prompt” and “existence of body problem”; she reported that she felt “disappointed” when seeing the image; she ranked the image as the worst among the five images generated from the same prompt.
Figure S2: An example of AU12 activation in response to a high-quality image generation. (a) The input text prompt. (b) The facial expression of the participant before seeing the generated image. (c) The generated image. (d) The facial expression of the participant after seeing the generated image (the frame with the highest estimated AU12 intensity). In the annotation survey, the participant gave ratings of 7 for each of image-text alignment, fidelity, and overall rating; he did not identify any issues with the image; he reported feeling “satisfied”, “surprised”, and “amused” when seeing the image; he ranked the image as the best among the five images generated from the same prompt.
B Data Collection
This section provides more details of the collection of the FERGI dataset not provided in Sec. 3. Figs. S3, S4, S5, S6, S7, S8 and S9 display screenshots of an example of one session in data collection.
B.1 Participants
39 participants [^1] The user IDs are up to 040 as shown in Tab. S1 because the same participant used both 027 and 028 due to some technical issues. Though, no data was collected under the user ID 027. were recruited from the SONA system of University of California San Diego (UCSD). and received academic credits as compensation. They underwent the informed consent process and received instruction for the study from the researcher via a videoconference and completed the study asynchronously afterwards on their personal computers. The videos of 3 participants failed to record and displayed only a logo of OBS Studio (the reason has not yet been fully understood by the researchers); 1 participant dropped out of study under the instruction of the researcher because the only available personal device was a tablet, which is not compatible with some procedures of the study; 1 participant did not complete any study sessions asynchronously at all after receiving the instruction in the videoconference. Among the other 34 participants with valid data and videos recorded, only 1 participant chose not to permit video recordings being shared with other researchers for research purposes; therefore, we decided to also exclude the data of this participant in our analysis for enhancing reproducibility of our study within the research community. We ended up with a dataset of 33 participants that will be made available to researchers for research purposes.
B.2 Prompt Creation
@@ B.2.1 Structured Input
After participants finish with the input sections and click “generate prompt”, a prompt will be generated based on all of the inputs with LLM (OpenAI gpt3.5-turbo model). Participants can freely adjust the generated prompt as the final prompt. Screenshots of an example of the inputs and prompt generation for a structured input session are shown in Figs. S3 and S4.
B.2.2 Free-Form Input
@@
B.3 Image Presentation
Screenshots of an example of the pages for image presentation and its baseline are shown in Figs. S6 and S7.
B.3.1 Image Annotation
@@ C AU Model Training
C.1 Datasets
The DISFA dataset [53] contains facial video recordings of 27 participants’ spontaneous facial expression while viewing video clips with approximately $130,000$ frames in total. Each frame is manually annotated by a human expert with intensities of AU1 (inner brow raiser), AU2 (outer brow raiser), AU4 (brow lowerer), AU5 (upper lid raiser), AU6 (cheek raiser), AU9 (nose wrinkler), AU12 (lip corner puller), AU15 (lip corner depressor), AU17 (chin raiser), AU20 (lip stretcher), AU25 (lips part), and AU26 (jaw drop) on a scale of 0 to 5. See Fig. S10 for a visual reference guide for the analyzed AUs.
Figure S10: A visual reference guide for the analyzed AUs extracted from [56].
@@ C.3 Model Training
For model training, we employ the Adam optimizer with an initial learning rate of $10^{-4}$ for parameters of the last layer and an $10^{-5}$ for other parameters, a weight decay of $5\times 10^{-4}$, and a batch size of 64. We train the model on all data from the DISFA and DISFA+ datasets [53, 51] for a total of 3 epochs using a single NVIDIA GeForce GTX 1080Ti 11G GPU for about 1 hour.
D Data Filtering